# LiDAR, GNSS and IMU Sensor Alignment through Dynamic Time Warping to Construct 3D City Maps

### v1

@@ 5.4.2 Loop-Based Centreline Comparison
Table 3: Summary statistics of centreline and intersection alignment metrics averaged over all test sequences. CI: Centroidal Offset (m); MOD: Mean Orientation Deviation (°); RMSE: Root Mean Square Error of centreline deviation (m); SD: Standard Deviation (m). Relative improvements are computed with respect to the LiDAR-only baseline.
@@
5.4.3 Intersection-Based Alignment Evaluation
The alignment of the intersections provides key insight into the local geometric consistency of the 3D city maps. We assess spatial accuracy by matching centroids extracted from skeletonized LiDAR maps with those from a georeferenced reference. Figure 20 shows the distribution of 51 intersections in Perth CBD. Figures 18 and 19 compare the results of the LiDAR-only method and the proposed framework. Green points indicate ground-truth centroids; red points show those extracted from reconstructed maps. The baseline shows clear misalignments, especially at complex or peripheral junctions due to accumulated drift. Our proposed method yields close centroid correspondence, indicating higher alignment accuracy. Alignment is quantified using four metrics: centroid offset (CO, in meters), mean orientation deviation (MOD, in degrees), local envelope error (LEE, maximum deviation within a 10 m radius), and standard deviation (SD) of centroid offset.
Table 2 summarizes results for 15 selected intersections, with percentage improvements shown in the right most columns. Table 3 summarizes average centerline and intersection alignment metrics across all test sequences. The proposed approach reduces centroid offset by 61.5%, orientation deviation by 55.1%, and LEE by 59.4% over the baseline. These improvements are most pronounced at challenging junctions and map boundaries. Accurate intersection alignment supports critical tasks such as trajectory planning, and robust localization. These results demonstrate the effectiveness of the proposed GNSS-constrained fusion framework.
6 Conclusion
@@
7 Acknowledgment
This work was funded by the University of Jeddah, Jeddah, Saudi Arabia, under grant No.(UJ-24-SUTU-1290). The authors, therefore, thank the University of Jeddah for its technical and financial support.The authors also extend their appreciation to Mangoes Mapping Pty Ltd for their in-kind contribution of data collection equipments, including LiDAR and GNSS systems.

### v2

@@ IV-D2 Loop-Based Centreline Comparison
TABLE III: Summary statistics of centreline and intersection alignment metrics averaged over all test sequences. CI: Centroidal Offset (m); MOD: Mean Orientation Deviation (°); RMSE: Root Mean Square Error of centreline deviation (m); SD: Standard Deviation (m). Relative improvements are computed with respect to the LiDAR-only baseline.
@@
IV-D3 Intersection-Based Alignment Evaluation
The alignment of the intersections provides key insight into the local geometric consistency of the 3D city maps. We assess spatial accuracy by matching centroids extracted from skeletonized LiDAR maps with those from a georeferenced reference. Figures 14 and 14 compare the results of the LiDAR-only method and the proposed framework. Green points indicate ground-truth centroids; red points show those extracted from reconstructed maps. The baseline shows clear misalignments, especially at complex or peripheral junctions due to accumulated drift. Our proposed method yields close centroid correspondence, indicating higher alignment accuracy. Alignment is quantified using four metrics: centroid offset (CO, in meters), mean orientation deviation (MOD, in degrees), local envelope error (LEE, maximum deviation within a 10 m radius), and standard deviation (SD) of centroid offset.
Table II summarizes results for 15 selected intersections, with percentage improvements shown in the right most columns. Table III summarizes average centerline and intersection alignment metrics across all test sequences. The proposed approach reduces centroid offset by 61.5%, orientation deviation by 55.1%, and LEE by 59.4% over the baseline. These improvements are most pronounced at challenging junctions and map boundaries. Accurate intersection alignment supports critical tasks such as trajectory planning, and robust localization. These results demonstrate the effectiveness of the proposed GNSS-constrained fusion framework.
V Conclusion
@@
Acknowledgment
This work was funded by the University of Jeddah, Jeddah, Saudi Arabia, under grant No.(UJ-24-SUTU-1290). The authors, therefore, thank the University of Jeddah for its technical and financial support. This project is a collaborative effort between the University of Jeddah and the University of Western Australia. The authors also extend their appreciation to Mangoes Mapping Pty Ltd for their in-kind contribution of data collection equipment, including LiDAR and GNSS systems.